Caroline just landed her dream job upon graduating from college. Her new job is in sales and requires a car, so Caroline has decided to buy herself a new car that will cost $16,831. She anticipates that she will drive the car for 4 years and then sell the car for $10,377 at the end of the 4th year. Her annual operating costs of the car are expected to be $2,992 in the first year and will increase by $273 in each of the following years. What is the annual worth of owning and operating the vehicle? Assume that Caroline's money is worth 6% per year. (note: round your answer to the nearest cent and do not include spaces, currency signs, or commas; hint: if your answer is negative, be sure to include a negative sign).
Answer #1

ANSWER:

I = 6%

Aw = cost of the car(a/p,i,n) + aoc + increase in aoc(a/g,i,n) + salvage value(a/f,i,n)

aw = -16,831(a/p,6%,4) - 2,992 - 273(a/g,6%,4) + 10,377(a/f,6%,4)

aw = -16,831 * 0.2886 - 2,992 - 273 * 1.427 + 10,377 * 0.2286

aw = -4,857.42 - 2,992 - 389.57 + 2,372.18

aw = -5,866.81

so the annual worth of owning and operating the vehicle is -$5,866.81